A car is accelerating with a coffee cup inside of it. Assume it accelerates from rest with constant acceleration during which time the cup is on the point of slipping (μ_s=0.3). What is the velocity of the car 5s later?

a) 14.7 m/s

b) 24.4 m/s

c) 38.75 m/s
Answer #1

Acceleration of the car is

a = usg =0.3*9.8 =2.94 m/s2

From Kinematic equation

V = Vo+at

Since car starts from rest Vo=0

V=0+2.94*5

V=14.7 m/s